OpenStudy (anonymous):

NEED HELP NOW!!! Point K is graphed on a number line at 3. Point L is 11 units away from point K on the number line. What could be the coordinates of point L? Choose ALL answers that are correct. A. –14 B. –8 C. 7 D. 14

OpenStudy (campbell_st):

well think of it this way 3 + 11 = 3 - 11 = both answers are 11 units from 3

OpenStudy (anonymous):

-8 and 14

OpenStudy (anonymous):

llamas

OpenStudy (anonymous):

@KraziiKai, Wtf?

OpenStudy (anonymous):

hey yo gurl wutz ur number ima 57 yr old male wanna cyberchat @OpiGeode

OpenStudy (anonymous):

WHAT? WHY? -.-

OpenStudy (anonymous):

@xXPuziDestroyer69Xx

OpenStudy (anonymous):

lol what ?